# Perturbative and Non-linear Analyses of Gravitational Turbulence in Spacetimes with Stable Light Rings

### v1

@@ Abstract
Some black hole mimickers, as well as black strings and other higher-dimensional spacetimes, exhibit stable light rings-regions where light or high-frequency gravitational waves can be trapped. In these regions, linear perturbations decay slowly, raising the possibility of nonlinear instability mechanisms. In this work, we study the cubic nonlinear wave equation as a proxy for Einstein’s equations, using a four-dimensional model geometry that allows stable trapping. By employing a perturbative approach, we show that the nonlinear wave equation on the sphere with dissipative terms captures several features of the full nonlinear problem. This framework allows us to confirm a previous conjecture: all higher-order energy norms grow for arbitrarily small initial fluctuation amplitudes. Additionally, we analyze the system’s mode spectrum at late times, revealing an inertial range dominated by a direct energy cascade. These findings further support the notion that spacetimes with stable light rings develop weak, high-frequency radiation hair, which will not generically lead to instabilities.
@@ I Introduction
There are many theoretical proposals for compact objects  [^1] We will refer generically to compact objects as those with compactness $\mathcal{C}=M/R\geq 1/3$, where $M$ and $R$ denote, respectively, the mass and the radius of the object or star., whose existence could violate the final state conjecture, with varying levels of development [11]. Probably the first time this idea appears in the literature is as a “geon,” introduced by Wheeler [12] searching for a self-gravitating configuration of the electromagnetic field. This is an example of a BH alternative that exists within a theory admitting BH solutions, such as General Relativity. In fact, although no stable geons were found, Wheeler’s idea survived in the form of boson stars [13, 14, 15], gravitational condensates of a scalar field. Other proposals involve mechanisms that prevent gravitational collapse from ever forming event horizons (typically involving quantum gravity phenomenology), such as fuzzballs [16, 17], gravastars [18], or anti-de-Sitter (AdS) black shells [19, 20, 21, 22, 23, 24]. The term BH mimickers is sometimes used to refer generically to both classes of objects, and we refer the reader to [25] for a comprehensive review on the topic. Here, we will focus on horizonless models, i.e., compact objects which do not have an event horizon, and are not singular either. We will briefly comment on the role of horizons later on. This class of BH mimickers was dubbed “strong” BH imposters in Ref. [11].
@@
GW observations from compact binary mergers provide a means to test the BH hypothesis against various BH mimicker models [26, 27, 28, 29, 30, 31, 32, 33]. This can shed light towards a possible ultraviolet completion of General Relativity [27], and help tackle the dark matter problem [34, 35]. Horizonless compact objects’ most stark signature is the presence of GW echoes [36, 37, 38, 39, 27, 25, 40]. No evidence for this has been found in current observations [33]. Gravitational-wave echo searches are typically based in parametrized models [41, 42, 43]. However, different BH mimicker predict GW echoes arriving with vastly different time delays and amplitudes, making it challenging to rule out certain models only based on current observations.
@@
An alternative approach is to investigate the generic dynamical properties of these objects. This implies questioning, e.g., how would such an object form dynamically, and its linear and nonlinear stability. Generically, horizonless objects that are also compact possess stable light rings (LRs). These are circular orbits where light rays can remain trapped indefinitely. Ref. [44] showed that the presence of stable LRs was associated to a slow decay of linear perturbations, and it was conjectured that this would lead to a nonlinear instability. The physical insight was that the trapping of high frequency modes for a very long time could trigger the formation of BHs, or the collapse or dispersion of the whole configuration. We refer to this as the LR instability conjecture. The goal of this work is to examine some of its fundamental aspects.
@@
Following the publication of [44], Ref. [45] investigated the perturbative properties of BH mimickers with stable LRs. In particular, as it was already known for some stellar models [46], these objects have long-lived modes. Heuristic arguments favoring the presence of a nonlinear instability were discussed, as well. This problem has also been investigated from the numerical point of view, focusing on particular classes of mimickers that are solutions of well posed equations of motion. Ref. [47] showed that radial perturbations of scalarized, charged BHs with a stable LR, do not trigger any instability. In the context of boson stars, Ref. [48] claimed a generic instability for rotating boson stars with stable LRs, with a moderate timescale. The fate of the instability depends on the model under consideration, in some cases resulting in the collapse to a BH, while in others the star migrates towards a less compact configuration, without LRs. However, Ref. [49] was able to track the relaxation of a remnant boson star found after the merger of two such objects, where the remnant had similar properties as Ref. [48], without any signs of an instability. Therefore, the numerical evidence on this topic is not conclusive enough, at the time of writing this article. Analytical insights into this problem prove, therefore, especially valuable [50].
@@
A related problem is the stability of higher dimensional black strings and black rings, which also can possess stable LRs. This was first studied in detail in [51], proving also slow decay of linear perturbations, and conjecturing that this could also trigger a nonlinear instability. These problems exhibit similarities with the turbulent instability of pure anti-de Sitter (AdS) space. This spacetime, with reflective boundary conditions, is generically unstable against scalar perturbations [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63]. Although there are certain classes of initial conditions for which there is no such instability [54, 64, 56, 65, 57], it is otherwise generic and needs only of two elements: (i) sufficient trapping of radiation, and (ii) a focusing mechanism, that localizes the radiation in a sufficiently small region close to the origin. The arguments in favor of the instability of BH mimickers with stable LRs invoke precisely this same two ingredients. Another system with similar features (rotating BHs in AdS) have also been shown numerically to be unstable against the development of high frequency hair [66].
@@
Our work is mainly motivated by a recent study [67], which investigates the dynamics of a defocusing cubic wave equation on a spacetime with stable trapping. By numerically evolving initial data localized within the stably trapped region, they identified a direct turbulent cascade in angular modes–a process in which higher and higher multipoles become excited, ultimately dominating the evolution. This mechanism arises due to the slow dispersion within the stable trapping region, where nonlinear interactions promote an energy transfer to increasingly small-scale structures. Crucially, based on their numerical results they conjectured that higher-order norms (or energies) of the solution can grow unbounded in time for generic initial data, in stark contrast with the decay properties typically seen when trapping is unstable (as in asymptotically flat black hole spacetimes like Schwarzschild). They further speculated that, even if the same turbulent behavior were to manifest for gravitational perturbations of compact objects or black string spacetimes in general relativity, it would not necessarily precipitate gravitational collapse or singularity formation.
@@
In this work, we will prove the aforementioned conjecture regarding the unbounded growth of higher-order norms in the perturbative regime. Our setup is the same as in Ref. [67], featuring in particular stable trapping, which underlies the slow dispersive properties driving the turbulent cascade. Establishing these results in a controlled, perturbative setting not only reinforces the numerical findings of Ref. [67], but also clarifies how stable trapping can fundamentally alter the expected asymptotic behavior of nonlinear waves. Additionally we show that, unlike in pure AdS (see, for instance, Refs. [52, 53, 56]), the confinement properties of this model are weak. Indeed, even at high frequencies, the radiation can be distributed in a number of overtones, which are long-lived and localized not at the LR itself, but within a “light shell” with non negligible thickness. Therefore, arguments based on the hoop conjecture to invoke the formation of BHs at the LR are bound to fail.
@@
The structure of the paper is as follows. First, we revisit the model spacetime and nonlinear wave equation of Ref. [67] in Sec. II. In Sec. III we solve the linear problem, studying the quasinormal mode spectrum of the system, and identify properties of the long-lived modes. We then examine the stability question from a mathematical point of view in Sec. IV. We compute the growth rate of higher-order norms within the perturbative regime for a simplified configuration. Furthermore, we show how this calculation accurately estimates the growth of the norms obtained from the numerical evolution of the nonlinear wave equation in a spacetime with stable trapping. Lastly, we examine the development of a turbulent regime, and the end state of this process, in Sec. V, and we conclude with a brief discussion of future prospects in Sec.VI. We relegate most of the technical details to Apps. A to D. In this work we use geometric units $G=1=c$, and choose the mostly plus signature of the metric ($+$)(-$+$)( - + + + ).
@@ II The Setup: A simple geometry with a stable light ring
While the spacetime we are considering is not a solution of vacuum Einstein’s equations, we motivate the choice of the function $f$ by requiring that it has a pair of LRs at tunable locations. A simple functional choice that achieves this behaviour is
@@
where $\gamma_{ij}$ is the spatial metric on the $t=\mathrm{const}$ slice labeled by $\Sigma_{t}$. If we turn off the nonlinearity, i.e., when $\kappa=0$, these nonlinear norms are bound uniformly, and the local energy decays. More precisely, if $\Omega$ is a non-empty and bounded compact region, which does not change with time, and we denote by $\norm{D_{\Omega}^{(k)}\Phi(t)}^{2}$, the norm defined in Eq. (12), but integrating only over the compact region, i.e., $\Sigma_{t}\cap\Omega$, we have [67]
@@ III Linear Perturbations
Let us solve the linear equation $\mathcal{O}_{\ell}\phi_{\ell}=0$ subject to out-going boundary conditions as $r\to\infty$, and look for solutions that are regular at the origin. This regularity condition can be seen as $\phi_{\ell}\sim r^{\ell+1}$ as $r\to 0$  [^2] Irregular solutions behave as $\phi_{\ell}\sim r^{-\ell}$.. Solutions satisfying these boundary conditions are quasinormal modes (QNMs). Using a shooting method we numerically solve the equation in the frequency domain by writing $\phi_{\ell}=e^{-i\omega_{\ell n}t}e_{\ell n}(r)$.
@@
This numerical method allows to obtain the fundamental QNM frequency for the first angular harmonics. The results can be visualized by writing $\omega_{\ell n}=\nu_{\ell n}-i/\tau_{\ell n}$, where $\nu_{\ell n}$ is the real part of the frequency of the QNM, and $\tau_{\ell n}$ its damping time. In particular, Fig. 2 shows how the damping time $\tau_{\ell n}R\sim e^{\gamma\ell}$ grows exponentially with $\ell$, with an exponential factor which can be computed directly from a WKB expansion (as shown in App. A), in the eikonal limit, as
@@
For the higher angular harmonics, Fig. 2 shows that these numerical solutions are in agreement with the WKB approximation, as well as a method based on Breit–Wigner resonances [46, 73]. The details of these approximate methods are presented in Appendices A and B, respectively. Notice that the direct integration method requires higher numerical precision to accurately resolve the higher $\ell$ harmonics. Since this becomes quickly a computationally expensive task, we only compute up to $\ell=10$.
Figure 2: Oscillation frequency (top) and damping time (bottom) of the first few QNMs for the modeled spacetime with $\alpha=0.166$. The exact values, computed with a direct integration, are shown as orange stars. Filled markers are computed using the WKB approximation, whereas the empty markers have been computed through the Breit–Wigner (BW) method. As expected, all three methods start showing good agreement for the higher harmonics $\ell\gtrsim 6$.
@@ IV Growth of the Norms
Let us now examine the stability of the nonlinear problem, i.e., when $\kappa\neq 0$. While the nonlinear energy norm (10) is conserved, the behavior of the higher order Sobolev norms (12) does not have to. Note that one can also can define higher order energy-like norms as
@@
These norms, as opposed to the Sobolev norms (12), involve time derivatives, as well as the potential term. Ref. [67] presented numerical evidence that second order derivatives (related to the second order Sobolev norms) grow within their simulation time for sufficiently large initial data. It was conjectured that all norms with $k>1$ grow for arbitrarily small initial data. In this Section we address this question by computing the rate at which both Sobolev and energy-like higher order norms grow, in the perturbative regime (therefore, for arbitrarily small initial data), in a simplified model. Lastly, we show, with additional numerical evidence, that the growth of the norms in the $3+1$ dimensional model follows a similar rate.
@@ IV.1 Two-Dimensional Model
The detailed solution is provided in the App. C. What matters for the current discussion is that the resonant terms on the right side of the above equations cause terms that grow secularly. To NLO we find linear growth in time, e.g., $a_{1}\propto t\sin(nt)$, and to NNLO we observe terms that grow like $t^{2}$. This means that the NLO solution is only valid up to times when $t\leq\mathscr{O}(\epsilon^{-1})$, and the NNLO solution is only valid up to $t\leq\mathscr{O}(\epsilon^{-1/2})$. A two-timescale expansion [74] would be necessary to find a solution valid for larger timescales.
We can now calculate the equivalents of the Sobolev and energy-like higher-order norms in this situation. While the computation is straightforward, it is somewhat lengthy, and further details can be found in App. C. In this case, the Sobolev and energy higher-order norms are expressed as follows:
@@
for any function $f(t)$, then, the rolling average of the Sobolev norm, after considering the time dependent mode coefficients, explicitly given in App. C, is
@@
It is now explicit that $\langle\norm{D^{(k)}\Phi}^{2}\rangle_{T}\supset\epsilon^{6}3^{2k}T^{2}$, whereas the other terms are either constants, or purely oscillatory terms. Thus, if we fix the period $T$ and the amplitude of the initial data $\epsilon$, we find that the higher-order norms have grown more relative to the lower-order norms. This is consistent with a direct energy cascade–if energy is flowing towards higher frequency modes, this must lead to faster growth of the higher energy norms since those weigh more heavily than the high-frequency content.
A similar behavior occurs for the energy-like higher-order norms. As shown in App. C, it leads to
@@
We have also checked that the NNLO term vanishes exactly, even before taking the average. However, for the higher-order norms, we observe the same behavior for both the Sobolev and energy-like norms, i.e., the norms grow as $\epsilon^{6}c_{k}T^{2}$, where $c_{k}$ is some monotone function of $k$, which depends on the precise definition of the norm. Note that the energy-like norms must have $c_{k=1}=0$, whereas this restriction does not apply to the Sobolev norms, which can grow even at first order.

### v2

# Perturbative and nonlinear analyses of gravitational turbulence in spacetimes with stable light rings

@@ Abstract
Some black hole mimickers, as well as black strings and other higher-dimensional spacetimes, exhibit stable light rings-regions where light or high-frequency gravitational waves can be trapped. In these regions, linear perturbations decay slowly, raising the possibility of nonlinear instability mechanisms. In this work, we study the cubic nonlinear wave equation as a proxy for Einstein’s equations, using a four-dimensional model geometry that allows stable trapping. By employing a perturbative approach, and neglecting the backreaction onto the spacetime, we show that the nonlinear wave equation on the sphere with dissipative terms captures several features of the full nonlinear problem. This framework allows us to confirm a previous conjecture: all higher-order energy norms grow for arbitrarily small initial fluctuation amplitudes. Additionally, we analyze the system’s mode spectrum at late times, revealing an inertial range dominated by a direct energy cascade. These findings further support the notion that spacetimes with stable light rings develop weak, high-frequency radiation hair, which will not generically lead to instabilities.
@@ I Introduction
There are many theoretical proposals for compact objects  [^1] We will refer generically to compact objects as those with compactness $\mathcal{C}=M/R\geq 1/3$, where $M$ and $R$ denote, respectively, the mass and the radius of the object or star., whose existence could violate the final state conjecture, with varying levels of development [11]. Probably the first time this idea appears in the literature is as a “geon,” introduced by Wheeler [12] searching for a self-gravitating configuration of the electromagnetic field. This is an example of a BH alternative that exists within a theory admitting BH solutions, such as General Relativity. In fact, although no stable geons were found, Wheeler’s idea survived in the form of boson stars [13, 14, 15], gravitational condensates of a scalar field. Other proposals involve mechanisms that prevent gravitational collapse from ever forming event horizons (typically involving quantum gravity phenomenology), such as fuzzballs [16, 17], gravastars [18], or anti–de Sitter (AdS) black shells [19, 20, 21, 22, 23, 24]. The term BH mimickers is sometimes used to refer generically to both classes of objects, and we refer the reader to [25] for a comprehensive review on the topic. Here, we will focus on horizonless models, i.e., compact objects which do not have an event horizon, and are not singular either. We will briefly comment on the role of horizons later on. This class of BH mimickers was dubbed “strong” BH imposters in Ref. [11].
@@
GW observations from compact binary mergers provide a means to test the BH hypothesis against various BH mimicker models [26, 27, 28, 29, 30, 31, 32, 33]. This can shed light toward a possible ultraviolet completion of General Relativity [27], and help tackle the dark matter problem [34, 35]. Horizonless compact objects’ most stark signature is the presence of GW echoes [36, 37, 38, 39, 27, 25, 40]. No evidence for this has been found in current observations [33]. Gravitational-wave echo searches are typically based in parametrized models [41, 42, 43]. However, different BH mimicker predict GW echoes arriving with vastly different time delays and amplitudes, making it challenging to rule out certain models only based on current observations.
@@
An alternative approach is to investigate the generic dynamical properties of these objects. This implies questioning, e.g., how would such an object form dynamically, and its linear and nonlinear stability. Generically, horizonless objects that are also compact possess stable light rings (LRs). These are circular orbits where light rays can remain trapped indefinitely. Reference [44] showed that the presence of stable LRs was associated to a slow decay of linear perturbations, and it was conjectured that this would lead to a nonlinear instability. The physical insight was that the trapping of high frequency modes for a very long time could trigger the formation of BHs, or the collapse or dispersion of the whole configuration. We refer to this as the LR instability conjecture. The goal of this work is to examine some of its fundamental aspects.
@@
Following the publication of [44], Ref. [45] investigated the perturbative properties of BH mimickers with stable LRs. In particular, as it was already known for some stellar models [46], these objects have long-lived modes. Heuristic arguments favoring the presence of a nonlinear instability were discussed, as well. This problem has also been investigated from the numerical point of view, focusing on particular classes of mimickers that are solutions of well-posed equations of motion. Ref. [47] showed that radial perturbations of scalarized, charged BHs with a stable LR, do not trigger any instability. In the context of boson stars, Ref. [48] claimed a generic instability for rotating boson stars with stable LRs, with a moderate timescale. The fate of the instability depends on the model under consideration, in some cases resulting in the collapse to a BH, while in others the star migrates toward a less compact configuration, without LRs. However, Ref. [49] was able to track the relaxation of a remnant boson star found after the merger of two such objects, where the remnant had similar properties as Ref. [48], without any signs of an instability. Therefore, further numerical research and analytical insights [50] on these systems are paramount.
@@
A related problem is the stability of higher dimensional black strings and black rings, which also can possess stable LRs. This was first studied in detail in [51], proving also slow decay of linear perturbations, and conjecturing that this could also trigger a nonlinear instability. These problems exhibit similarities with the turbulent instability of pure anti–de Sitter (AdS) space. This spacetime, with reflective boundary conditions, is generically unstable against scalar perturbations [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64]. Although there are certain classes of initial conditions for which there is no such instability [54, 65, 56, 66, 57], it is otherwise generic and needs only of two elements: (i) sufficient trapping of radiation, and (ii) a focusing mechanism, that localizes the radiation in a sufficiently small region close to the origin. The arguments in favor of the instability of BH mimickers with stable LRs invoke precisely this same two ingredients. Another system with similar features (rotating BHs in AdS) have also been shown numerically to be unstable against the development of high frequency hair [67].
@@
Our work is mainly motivated by a recent study [68], which investigates the dynamics of a defocusing cubic wave equation on a spacetime with stable trapping. By numerically evolving initial data localized within the stably trapped region, they identified a direct turbulent cascade in angular modes—a process in which higher and higher multipoles become excited, ultimately dominating the evolution. This mechanism arises due to the slow dispersion within the stable trapping region, where nonlinear interactions promote an energy transfer to increasingly small-scale structures. Crucially, based on their numerical results they conjectured that higher-order norms (or energies) of the solution can grow unbounded in time for generic initial data, in stark contrast with the decay properties typically seen when trapping is unstable (as in asymptotically flat black hole spacetimes like Schwarzschild). They further speculated that, even if the same turbulent behavior were to manifest for gravitational perturbations of compact objects or black string spacetimes in general relativity, it would not necessarily precipitate gravitational collapse or singularity formation.
@@
In this work, we will prove the aforementioned conjecture regarding the unbounded growth of higher-order norms in the perturbative regime. Our setup is the same as in Ref. [68], featuring in particular stable trapping, which underlies the slow dispersive properties driving the turbulent cascade. In this model, since backreaction on the spacetime is not included, the nonlinear behavior we study arises purely from the self-interaction of the scalar field. Establishing these results in a controlled, perturbative setting not only reinforces the numerical findings of Ref. [68], but also clarifies how stable trapping can fundamentally alter the expected asymptotic behavior of nonlinear waves. Additionally we show that, unlike in pure AdS (see, for instance, Refs. [52, 53, 56]), the confinement properties of this model are weak. Indeed, even at high frequencies, the radiation can be distributed in a number of overtones, which are long-lived and localized not at the LR itself, but within a “light shell” with non negligible thickness. Therefore, arguments based on the hoop conjecture to invoke the formation of BHs at the LR are bound to fail.
@@
The structure of the paper is as follows. First, we revisit the model spacetime and nonlinear wave equation of Ref. [68] in Sec. II. In Sec. III we solve the linear problem, studying the quasinormal mode spectrum of the system, and identify properties of the long-lived modes. We then examine the stability question from a mathematical point of view in Sec. IV. We compute the growth rate of higher-order norms within the perturbative regime for a simplified configuration. Furthermore, we show how this calculation accurately estimates the growth of the norms obtained from the numerical evolution of the nonlinear wave equation in a spacetime with stable trapping. Lastly, we examine the development of a turbulent regime, and the end state of this process, in Sec. V, and we conclude with a brief discussion of future prospects in Sec.VI. We relegate most of the technical details to Appendices A to D. In this work we use geometric units $G=1=c$, and choose the mostly plus signature of the metric ($+$)(-$+$)( - + + + ).
@@ II The Setup: A simple geometry with a stable light ring
While the spacetime we are considering is not a solution of vacuum Einstein’s equations, we motivate the choice of the function $f$ by requiring that it has a pair of LRs at tunable locations. A simple functional choice that achieves this behavior is
@@
where $\gamma_{ij}$ is the spatial metric on the $t=\mathrm{const}$ slice labeled by $\Sigma_{t}$. If we turn off the nonlinearity, i.e., when $\kappa=0$, these nonlinear norms are bound uniformly, and the local energy decays. More precisely, if $\Omega$ is a nonempty and bounded compact region, which does not change with time, and we denote by $\norm{D_{\Omega}^{(k)}\Phi(t)}^{2}$, the norm defined in Eq. (12), but integrating only over the compact region, i.e., $\Sigma_{t}\cap\Omega$, we have [68]
@@ III Linear Perturbations
Let us solve the linear equation $\mathcal{O}_{\ell}\phi_{\ell}=0$ subject to outgoing boundary conditions as $r\to\infty$, and look for solutions that are regular at the origin. This regularity condition can be seen as $\phi_{\ell}\sim r^{\ell+1}$ as $r\to 0$  [^2] Irregular solutions behave as $\phi_{\ell}\sim r^{-\ell}$.. Solutions satisfying these boundary conditions are quasinormal modes (QNMs). Using a shooting method we numerically solve the equation in the frequency domain by writing $\phi_{\ell}=e^{-i\omega_{\ell n}t}e_{\ell n}(r)$.
@@
This numerical method allows to obtain the fundamental QNM frequency for the first angular harmonics. The results can be visualized by writing $\omega_{\ell n}=\nu_{\ell n}-i/\tau_{\ell n}$, where $\nu_{\ell n}$ is the real part of the frequency of the QNM, and $\tau_{\ell n}$ its damping time. In particular, Fig. 2 shows how the damping time $\tau_{\ell n}R\sim e^{\gamma\ell}$ grows exponentially with $\ell$, with an exponential factor which can be computed directly from a Wentzel–Kramers–Brillouin (WKB) expansion (as shown in Appendix A), in the eikonal limit, as
@@
For the higher angular harmonics, Fig. 2 shows that these numerical solutions are in agreement with the WKB approximation, as well as a method based on Breit-Wigner resonances [46, 74]. The details of these approximate methods are presented in Appendices A and B, respectively. Notice that the direct integration method requires higher numerical precision to accurately resolve the higher $\ell$ harmonics. Since this becomes quickly a computationally expensive task, we only compute up to $\ell=10$.
Figure 2: Oscillation frequency (top) and damping time (bottom) of the first few QNMs for the modeled spacetime with $\alpha=0.166$. The exact values, computed with a direct integration, are shown as orange stars. Filled markers are computed using the WKB approximation, whereas the empty markers have been computed through the Breit-Wigner (BW) method. As expected, all three methods start showing good agreement for the higher harmonics $\ell\gtrsim 6$.
@@ IV Growth of the Norms
Let us now examine the stability of the nonlinear problem, i.e., when $\kappa\neq 0$. While the nonlinear energy norm (10) is conserved, the behavior of the higher order Sobolev norms (12) does not have to. Note that one can also can define higher order energylike norms as
@@
These norms, as opposed to the Sobolev norms (12), involve time derivatives, as well as the potential term. Ref. [68] presented numerical evidence that second order derivatives (related to the second order Sobolev norms) grow within their simulation time for sufficiently large initial data. It was conjectured that all norms with $k>1$ grow for arbitrarily small initial data. In this Section we address this question by computing the rate at which both Sobolev and energylike higher order norms grow, in the perturbative regime (therefore, for arbitrarily small initial data), in a simplified model. Lastly, we show, with additional numerical evidence, that the growth of the norms in the $3+1$ dimensional model follows a similar rate.
@@ IV.1 Two-dimensional Model
The detailed solution is provided in the Appendix C. What matters for the current discussion is that the resonant terms on the right side of the above equations cause terms that grow secularly. To NLO we find linear growth in time, e.g., $a_{1}\propto t\sin(nt)$, and to NNLO we observe terms that grow like $t^{2}$. This means that the NLO solution is only valid up to times when $t\leq\mathscr{O}(\epsilon^{-1})$, and the NNLO solution is only valid up to $t\leq\mathscr{O}(\epsilon^{-1/2})$. A two-timescale expansion [75] would be necessary to find a solution valid for larger timescales.
We can now calculate the equivalents of the Sobolev and energylike higher-order norms in this situation. While the computation is straightforward, it is somewhat lengthy, and further details can be found in Appendix C. In this case, the Sobolev and energy higher-order norms are expressed as follows:
@@
for any function $f(t)$, then, the rolling average of the Sobolev norm, after considering the time dependent mode coefficients, explicitly given in Appendix C, is
@@
It is now explicit that $\langle\norm{D^{(k)}\Phi}^{2}\rangle_{T}\supset\epsilon^{6}3^{2k}T^{2}$, whereas the other terms are either constants, or purely oscillatory terms. Thus, if we fix the period $T$ and the amplitude of the initial data $\epsilon$, we find that the higher-order norms have grown more relative to the lower-order norms. This is consistent with a direct energy cascade—if energy is flowing toward higher frequency modes, this must lead to faster growth of the higher energy norms since those weigh more heavily than the high-frequency content.
A similar behavior occurs for the energylike higher-order norms. As shown in Appendix C, it leads to
@@
We have also checked that the NNLO term vanishes exactly, even before taking the average. However, for the higher-order norms, we observe the same behavior for both the Sobolev and energylike norms, i.e., the norms grow as $\epsilon^{6}c_{k}T^{2}$, where $c_{k}$ is some monotone function of $k$, which depends on the precise definition of the norm. Note that the energylike norms must have $c_{k=1}=0$, whereas this restriction does not apply to the Sobolev norms, which can grow even at first order.